# Reconstruction of Caterpillar Tanglegrams

### v1

@@ 4. Reconstruction of caterpillar tanglegrams
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a tanglegram of size $n\geq 5$, where the left tree has distance-to-root labeling $\{v_{1},\ldots,v_{n}\}$. If $\sigma$ matches $v_{1}$ with a leaf of the cherry of the right tree, then $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
@@
The case $n=5$ can be checked directly, so throughout, we assume $n\geq 6$.
We first prove that $\mathcal{D}^{m}(\mathcal{T})$ determines whether or not $v_{1}$ is matched to a leaf of the cherry of the right tree. Let $\{u_{1},\ldots,u_{n}\}$ be the distance-to-root labeling of the right tree. If $v_{1}$ is adjacent to a leaf of the cherry, we may assume without loss of generality that the distance-to-root labeling is such that $v_{1}$ is adjacent to $u_{n}$. Observe that if $v_{1}$ is matched to $u_{n}$, then at least $n-1$ cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to a leaf in the cherry of $R^{\prime}$. If $v_{1}$ is matched to $u_{n-2}$, then between two and three cards (corresponding to the removal of $u_{n-1},u_{n}$, and possibly $v_{1}$) have $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. If $v_{1}$ is matched to some $u_{i}$ with $i\leq n-3$, then at most one card has $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. Hence, by counting cards in which $v_{1}^{\prime}$ is matched to the cherry of $R^{\prime}$, we can determine from $\mathcal{D}^{m}(\mathcal{T})$ if $v_{1}$ is matched to $u_{n}$.
We now show that when $v_{1}$ is matched to $u_{n}$, $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$. There are various cases and subcases based on which vertices $v_{2}$ and potentially $v_{3}$ are matched to in $\mathcal{T}$. These are collected in Table 1. Observe that we can distinguish the case where $v_{2}$ is matched to $u_{j}$ with $j<n-2$, as there exists a unique card, $\mathcal{T}-v_{1}u_{n}$, where $v_{1}^{\prime}$ is not matched to the cherry of $R^{\prime}$. Using this card also allows us to reconstruct $\mathcal{T}$.
Table 1. Case analysis for $v_{1}u_{n}\in\sigma$. We assume without loss of generality that in the card $\mathcal{T}^{\prime}=(L^{\prime},R^{\prime},\sigma^{\prime})$, the distance-to-root labeling of $R^{\prime}$ is chosen so that $v_{1}^{\prime}$ is matched to $u_{n-1}^{\prime}$ whenever possible.
@@
For the remaining four subcases, observe that the number of cards where $v_{1}^{\prime}$ and $v_{2}^{\prime}$ are matched to the cherry of $R^{\prime}$ is $n$ (if $v_{2}$ is matched to $u_{n-1}$ and $v_{3}$ is matched to $u_{n-2}$), $n-2$ (if $v_{2}$ is matched to $u_{n-1}$ but $v_{3}$ is not matched to $u_{n-2}$), $3$ (if $v_{2}$ is matched to $u_{n-2}$ and $v_{3}$ is matched to $u_{n-1}$), or $1$ (if $v_{2}$ is matched to $u_{n-2}$ but $v_{3}$ is not matched to $u_{n-1}$). As $n\geq 6$, these are all distinct, so $\mathcal{D}^{m}(\mathcal{T})$ allows us to distinguish these four subcases. It now suffices to show that we can reconstruct $\mathcal{T}$ in each of them.
@@
If $m=n$, then all cards are identical. If $m<n$, then the $m$ cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$ are identical with $v_{1}^{\prime},\ldots,v_{m-1}^{\prime}$ satisfying $v_{i}^{\prime}$ being matched to $u_{n-i}^{\prime}$, while $v_{m}^{\prime}$ is not matched to $u_{n-m}^{\prime}$. The other $n-m$ cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{i}^{\prime}$ matched to $u_{n-i^{\prime}}$ for $1\leq i\leq m$. Thus, we can determine $m$, and using one of the identical cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$, we can reconstruct $\mathcal{T}$.
@@
Finally, in the case where $v_{2}u_{n-2}\in\sigma$ and $v_{3}u_{n-1}\notin\sigma$, we have that $u_{n-1}$ is matched to some $v_{k}$ with $k>3$. There is a unique card, $\mathcal{T}-v_{k}u_{n-1}$, in which the cherry of $R^{\prime}$ is matched to both $v_{1}^{\prime}$ and $v_{2}^{\prime}$. Among the remaining cards, there are $n-k$ cards (corresponding to the removal of $v_{i}$ with $i>k$) where $v_{k}^{\prime}$ is matched to the cherry of $R^{\prime}$ and $k-1$ cards (corresponding to the removal of $v_{i}$ with $i<k$) where $v_{k-1}^{\prime}$ is matched to the cherry of $R^{\prime}$. Hence, we can identify $k$, and using the card $\mathcal{T}-v_{k}u_{n-1}$ allows us to reconstruct $\mathcal{T}$. ∎
@@
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a tanglegram of size $n\geq 5$. Then $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
@@
Let $\{v_{1},\ldots,v_{n}\}$ be the distance-to-root labeling of the left tree $L=C_{n}$ and let $\{u_{1},\ldots,u_{n}\}$ be the distance-to-root labeling of the right tree $R=C_{n}$. Furthermore, for a card $(L^{\prime},R^{\prime},\sigma^{\prime})\in\mathcal{D}^{m}(\mathcal{T})$, let $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ and $\{u_{1}^{\prime},\ldots,u_{n-1}^{\prime}\}$ denote the distance-to-root labelings of the induced subtrees $L^{\prime}$ and $R^{\prime}$.
Assume $v_{1}$ is matched to $u_{i}$ and $v_{2}$ is matched to $u_{j}$ with $i\neq j$. By Lemma 4.2, $\mathcal{T}$ is reconstructable if $i\geq n-1$, so we will assume $i<n-1$. Additionally, if $j\in\{n-1,n\}$, we set $j=n-1$.
If $i=1$, then at least $n-1$ of the cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to $u_{1}^{\prime}$. If instead $i\neq 1$, then there are at most two such cards (corresponding to $j=1$ or $i=2$). Thus, from $\mathcal{D}^{m}(\mathcal{T})$, we can determine whether or not $i=1$.
Suppose $i=1$. Let $v_{1},\ldots,v_{m}$ be a maximal sequence of vertices satisfying that $v_{i}$ is matched to $u_{i}$ for all $1\leq i\leq m$. If $m=n$, all cards are identical and the tanglegram is reconstructable. Since $m=n-1$ is not possible, we consider $m<n-1$. Then, in exactly $m$ identical cards (corresponding to $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$), we have $v_{k}^{\prime}$ matched to $u_{k}^{\prime}$ for $11\leq k\leq m-$ 1 and $v_{m}^{\prime}$ not matched to $u_{m}^{\prime}$. In the remaining cards, $v_{k}^{\prime}$ and $u_{k}^{\prime}$ are matched for $1\leq k\leq m$. Thus, we can identify $m$, and the $m$ identical cards resulting from $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$ allow us to reconstruct $\mathcal{T}$.
Suppose instead $i\neq 1$. In the cards of $\mathcal{D}^{m}(\mathcal{T})$, we consider which vertex is matched to $v_{1}^{\prime}$. From the cards other than $\mathcal{T}-v_{1}u_{i}$, we get $i-1$ cards with $v_{1}^{\prime}$ matched to $u_{i-1}^{\prime}$ (corresponding to the removal of $u_{k}$ with $1\leq k<i$) and $n-i$ cards with $v_{1}^{\prime}$ matched to $u_{i}^{\prime}$ (corresponding to the removal of $u_{k}$ with $i<k\leq n$). Which vertex $v_{1}^{\prime}$ is matched to on the remaining card, $\mathcal{T}-v_{1}u_{i}$, depends on the relationship of $i$ and $j$. The vertices matched to $v_{1}^{\prime}$ and their multiplicities in $\mathcal{D}^{m}(\mathcal{T})$ are collected in Table 2.
Table 2. For $21<i<n-$ 2, the vertices matched to $v_{1}^{\prime}$ in $\mathcal{D}^{m}(T)$ with multiplicities.
$j>i+1$ $j=i+1$ $j=i-1$ $j<i-1$
@@
$u_{i-1}^{\prime}\times(i-1)$ $u_{j}^{\prime}\times 1$
From Table 2, if $v_{1}^{\prime}$ is matched to exactly two $u_{k}^{\prime}$, then $j=i+1$ or $j=i-1$ and the multiplicities allow us distinguish between these cases. If $j=i+1$, then there is a maximal sequence of vertices $v_{1},\ldots,v_{m}$ satisfying that $v_{k}$ is matched to $u_{i+k-1}$ for $11\leq k\leq m-$ 1. Hence, there are $m$ identical cards with $v_{k}^{\prime}$ matched to $u_{i+k-1}^{\prime}$ for $11\leq k\leq m-$ 1 and $v_{m}^{\prime}$ not matched to $u_{i+m-1}^{\prime}$. However, in all other cards, $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ are matched to a set of $u$’s that are consecutively and increasingly indexed. Thus, we can identify $m$, and one of the $m$ identical cards resulting from removing $v_{k}u_{i+k-1}$ with $1\leq k\leq m$ allows us to reconstruct $\mathcal{T}$. If instead $j=i-1$, then a similar argument with $v_{k}u_{i-k+1}\in\sigma$ for $1\leq k\leq m$ shows that $\mathcal{T}$ is reconstructable.
Finally, we consider when $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ for three different values of $k$. If these indices are non-consecutive, then we can easily identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$. In this case, $\mathcal{T}$ is reconstructable from this card. If the indices are consecutive, then $j=i+2$ or $j=i-2$ and we consider the multiplicity of the cards in which $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ with $k$ maximum. In the former case, there is one such card and in the latter case, there are $n-i\geq 2$ such cards (since $i\leq n-1$). Hence, we can distinguish between these cases, identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$, and reconstruct $\mathcal{T}$. ∎
We now consider the remaining cases where exactly one of $L$ or $R$ is a caterpillar. Without loss of generality, $L=C_{n}$. We first consider the case that $R$ is type 1.

### v2

@@ 3. Types of trees
Let $T$ be a tree of type 1 with stripping $(i,Q)$, Then we have:
  1. (i)
    If $v$ is a strippable leaf, then TvT-v has stripping (i1,Q)(i-1,Q).
  2. (ii)
    If $v$ is a leaf of $Q$, then TvT-v has stripping $(i+j,Q^{\prime})$, where $(j,Q^{\prime})$ is the stripping of QvQ-v.
In particular, cards in $\mathcal{D}(T)$ have at least i1i-1 strippable leaves, and at least one card in $\mathcal{D}(T)$ has exactly i1i-1 strippable leaves; consequently, $i$ can be determined from $\mathcal{D}(T)$.
Let $T^{\prime}\in\mathcal{D}(T)$ with stripping $(j,Q^{\prime})$. Then we have one of the following:
  1. (a)
    j=i1j=i-1 and $Q^{\prime}=Q$;
  2. (b)
    $j=i$, $Q^{\prime}\in\mathcal{D}(Q)$, and $T^{\prime}=T-v$ for some leaf $v$ in a pending subtree of $Q$ that is of size at least $3$;
  3. (c)
    $j>i$, $Q=C_{2}\oplus(C_{1}\oplus)^{j-i-1}Q^{\prime}$, and $v$ is a leaf of the pending subtree $C_{2}$ of $Q$.
In particular, we can determine which leaves of $T^{\prime}$ correspond to leaves of $Q$ and which leaves correspond to strippable leaves of $T$. Moreover, we can determine which leaves of $T$ correspond to which pending subtree of $Q$, unless $Q$ is of type 2a or $Q=Q_{1}\oplus Q_{2}$ is of type 2b and $T^{\prime}=C_{1}\oplus(Q_{j}\oplus Q_{j})$ where $\left\langle Q_{j}\right\rangle>\left\langle Q_{3-j}\right\rangle$.
Lemma 3.6.
Let $T=T_{1}\oplus T_{2}$ be a tree of type 2, let $v$ be any leaf of $T$, and let $T^{\prime}=T-v$. If $T$ is not type 2a, we can determine a unique $j\in\{1,2\}$ such that $v\in T_{j}$ and $T^{\prime}=(T_{j}-v)\oplus T_{3-j}$. In particular, we can determine whether $v\in T_{1}$ or $v\in T_{2}$. If, in addition, $T^{\prime}$ is not type 2a, we can identify which pending subtree of $T^{\prime}$ corresponds to $T_{3-j}$.
4. Tanglegrams of small size
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$\mathcal{T}_{1}$$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$u_{1}$$u_{2}$$u_{3}$$u_{4}$$u_{5}$$\mathcal{T}_{2}$
$\mathcal{T}_{1}-v_{3}u_{2}=\mathcal{T}_{1}-v_{4}u_{1}=\mathcal{T}_{1}-v_{5}u_{3}$$\mathcal{T}_{2}-v_{1}u_{4}=\mathcal{T}_{2}-v_{2}u_{3}=\mathcal{T}_{2}-v_{3}u_{5}$$\mathcal{T}_{1}-v_{1}u_{4}=\mathcal{T}_{1}-v_{2}u_{5}$$\mathcal{T}_{2}-v_{4}u_{1}=\mathcal{T}_{2}-v_{5}u_{3}$
Figure 5. Two size $5$ catergrams that have the same multideck, and the cards in their deck. A distance-to-root labeling of the left and right trees of size $5$ catergrams are given as well as all representations of the cards in the deck.
Consider size $n$ rooted binary trees: When $n\leq 3$, the only rooted binary tree is the caterpillar $C_{n}$, for $n=4$, we have two trees, $C_{4}$ and $C_{2}\oplus C_{2}$, and for $n=5$, we have three trees, $C_{5}$, $C_{1}\oplus(C_{2}\oplus C_{2})$, and $C_{2}\oplus C_{3}$. The two binary trees of size $4$ have the same multidecks: $4$ copies of $C_{3}$.
Following the terminology of [5], we refer to tanglegrams of the form $(C_{n},C_{n},\sigma)$, where the left and right trees are both caterpillars, as catergrams. Clearly, the deck of a catergram only contains catergrams (and for $n\leq 3$, every size $n$ tanglegram is a catergram). If we denote the number of size $n$ catergrams by $c_{n}$, then $c_{1}=1$, and, as it is implicit in [5], for $n\geq 2$, we have $c_{n}=(n-2)!+\frac{n!-2(n-2)!}{4}=(n-2)!\frac{n^{2}-n+2}{4}$. In particular, $c_{2}=1$, $c_{3}=2$, and $c_{4}=7$. Moreover, every catergram of size $n$ (and for $n=4$ every tanglegram) has a deck that contains only catergrams of size n1n-1. Thus, the multideck of a size $n$ catergram can be described by the multiplicities of the size (n1)(n-1) catergrams in it (including 0). Therefore, there are at most $\binom{n+c_{n}-1}{c_{n}-1}$ possible decks for size $n$ catergrams. There are $17$ size $4$ tanglegrams, $7$ of these are catergrams, and there are at most $\binom{4+c_{3}-1}{c_{3}-1}=5$ multidecks for size $4$ tanglegrams. Consequently, even size $4$ catergrams are not reconstructable from their multideck.
Thus, the best we can hope for is reconstructing tanglegrams of size at least $5$, and by Lemma 2.17 we can reconstruct their left and right trees from their multideck. However, we find that even catergrams of size $5$ are not reconstructable from their multidecks, as the example in Figure 5 shows. Therefore the lower bound $6$ in Theorem 1.1 is best possible.
5. Reconstruction of catergrams
Throughout the remaining three sections, we freely use the fact that if $\mathcal{T}=(L,R,\sigma)$ is a tanglegram of size $n\geq 6$, the deck $\mathcal{D}(\mathcal{T})$ uniquely determines the trees $L$ and $R$. Consequently, as long as $\left\langle\mathcal{T}\right\rangle\geq 6$, we may freely assume that we know $L$ and $R$, and only need to determine $\sigma$ from $\mathcal{D}^{m}(\mathcal{T})$ (see Lemma 2.17). Also, without loss of generality, we will assume that $L=C_{n}$ and use the distance-to-root labeling $\{v_{1},\ldots,v_{n}\}$. Moreover, for cards $(C_{n-1},R^{\prime},\sigma^{\prime})$ we use subscripted symbols $\{v_{1}^{\prime},\ldots,v_{n-1}^{\prime}\}$ for the distance-to-root labeling of the left subtree.
In this section, we prove reconstruction for catergrams of size at least $6$. In this case, we use the distance-to-root labeling $\{u_{1},\ldots,u_{n}\}$ for the right subtree $R=C_{n}$, and for cards $T^{\prime}=(C_{n-1},C_{n-1},\sigma^{\prime})$, corresponding subscripted symbols $\{u_{1}^{\prime},\ldots,u_{n-1}^{\prime}\}$ for the distance-to-root labeling of the right subtree.
We first show that the multideck determines whether $v_{1}$ is matched to the cherry of $R$ or not.
Lemma 5.1.
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a catergram of size $n\geq 6$. We can determine from $\mathcal{D}^{m}(\mathcal{T})$ whether $v_{1}$ is matched to one of $u_{n},u_{n-1}$ or not.
@@ 5. Reconstruction of catergrams
If $v_{1}$ is adjacent to one of $u_{n},u_{n-1}$, we may assume without loss of generality that the distance-to-root labeling is such that $v_{1}$ is adjacent to $u_{n}$. Observe that if $v_{1}$ is matched to $u_{n}$, then at least $n-1\geq 5$ cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to a leaf in the cherry of $R^{\prime}$. If $v_{1}$ is matched to $u_{n-2}$, then precisely two or three cards (corresponding to the removal of $u_{n-1},u_{n}$, and possibly $v_{1}$) have $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. If $v_{1}$ is matched to some $u_{i}$ with $i\leq n-3$, then at most one card (namely $\mathcal{T}-v_{1}$) has $v_{1}^{\prime}$ matched to the cherry of $R^{\prime}$. Hence, by counting cards in which $v_{1}^{\prime}$ is matched to the cherry of $R^{\prime}$, we can determine from $\mathcal{D}^{m}(\mathcal{T})$ if $v_{1}$ is matched to one of $u_{n},u_{n-1}$ or not. ∎
We now handle the case in which $v_{1}$ is matched to the cherry of $R$.
Lemma 5.2.
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a catergram of size $n\geq 6$. If $\sigma$ matches $v_{1}$ with a leaf of the cherry of the right tree, then $\mathcal{T}$ is reconstructable from the multideck $\mathcal{D}^{m}(\mathcal{T})$.
Proof.
There are various cases and subcases based on which vertices $v_{2}$ and potentially $v_{3}$ are matched to in $\mathcal{T}$. These are collected in Table 1.
Observe that using $\mathcal{D}^{m}(\mathcal{T})$, we can distinguish the case where $v_{2}$ is matched to $u_{j}$ with j<n2j<n-2, as in this case, there exists a unique card, $\mathcal{T}-v_{1}u_{n}$, where $v_{1}^{\prime}$ is not matched to the cherry of $R^{\prime}$, whereas if $v_{2}$ is matched to one of $u_{n-2},u_{n-1}$, then $v_{1}^{\prime}$ is matched to the cherry of $R^{\prime}$ in all cards $\mathcal{T}^{\prime}$ of $\mathcal{D}^{m}(\mathcal{T})$.
If $v_{2}$ is matched to $u_{j}$ for some j<n2j<n-2, the card $\mathcal{T}-v_{1}u_{n}$ allows us to reconstruct $\mathcal{T}$.
Table 1. Case analysis for $v_{1}u_{n}\in\sigma$. We assume without loss of generality that in the card $\mathcal{T}^{\prime}=(L^{\prime},R^{\prime},\sigma^{\prime})$, the distance-to-root labeling of $R^{\prime}$ is chosen so that $v_{1}^{\prime}$ is matched to $u_{n-1}^{\prime}$ whenever it is matched to the cherry of $R^{\prime}$.
@@
Therefore we now assume that $v_{2}$ is matched to either $u_{n-1}$ or $u_{n-2}$. For the remaining four subcases, observe that the number of cards where $v_{1}^{\prime}$ and $v_{2}^{\prime}$ are matched to the cherry of $R^{\prime}$ is either $n$ (if $v_{2}$ is matched to $u_{n-1}$ and $v_{3}$ is matched to $u_{n-2}$), n2n-2 (if $v_{2}$ is matched to $u_{n-1}$ but $v_{3}$ is not matched to $u_{n-2}$), $3$ (if $v_{2}$ is matched to $u_{n-2}$ and $v_{3}$ is matched to $u_{n-1}$), or $1$ (if $v_{2}$ is matched to $u_{n-2}$ but $v_{3}$ is not matched to $u_{n-1}$). As $n\geq 6$, these are all distinct, so $\mathcal{D}^{m}(\mathcal{T})$ allows us to distinguish these four subcases. It now suffices to show that we can reconstruct $\mathcal{T}$ in each of them.
@@
If $m=n$, then all cards are identical. If $m<n$, then the $m$ cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$ are identical with $v_{1}^{\prime},\ldots,v_{m-1}^{\prime}$ satisfying $v_{i}^{\prime}$ being matched to $u_{n-i}^{\prime}$, while $v_{m}^{\prime}$ is not matched to $u_{n-m}^{\prime}$. The other nmn-m cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{i}^{\prime}$ matched to $u_{n-i}^{\prime}$ for $1\leq i\leq m$. Thus, we can determine $m$, and using one of the identical cards $\mathcal{T}-v_{i}u_{n-i+1}$ with $1\leq i\leq m$, we can reconstruct $\mathcal{T}$.
@@
Finally, in the case where $v_{2}u_{n-2}\in\sigma$ and $v_{3}u_{n-1}\notin\sigma$, we have that $u_{n-1}$ is matched to some $v_{k}$ with $k>3$. There is a unique card, $\mathcal{T}-v_{k}u_{n-1}$, in which the cherry of $R^{\prime}$ is matched to both $v_{1}^{\prime}$ and $v_{2}^{\prime}$. Among the remaining cards, there are nkn-k cards (corresponding to $\mathcal{T}-v_{i}$ with $i>k$) where $v_{k}^{\prime}$ is matched to the cherry of $R^{\prime}$ and k1k-1 cards (corresponding to $\mathcal{T}-v_{i}$ with $i<k$) where $v_{k-1}^{\prime}$ is matched to the cherry of $R^{\prime}$. Hence, we can identify $k$, and using the card $\mathcal{T}-v_{k}u_{n-1}$ allows us to reconstruct $\mathcal{T}$. ∎
@@
Let $\mathcal{T}=(C_{n},C_{n},\sigma)$ be a catergram of size $n\geq 6$. Then $\mathcal{T}$ is reconstructable from $\mathcal{D}^{m}(\mathcal{T})$.
@@
First note that Lemma 2.17 implies that the class of catergrams of size at least $6$ are deck-decidable.
Assume $v_{1}$ is matched to $u_{i}$ and $v_{2}$ is matched to $u_{j}$ with $i\neq j$. By Lemma 5.1, we can determine from $\mathcal{D}^{m}(\mathcal{T})$ whether $i\geq n-1$ or i<n1i<n-1. By Lemma 5.2, $\mathcal{T}$ is reconstructable if $i\geq n-1$, so we will assume i<n1i<n-1. Additionally, if $j\in\{n-1,n\}$, we set j=n1j=n-1.
If $i=1$, then at least n1n-1 of the cards in $\mathcal{D}^{m}(\mathcal{T})$ have $v_{1}^{\prime}$ matched to $u_{1}^{\prime}$. If instead $i\neq 1$, then there are at most two such cards (if $i=2$, then $\mathcal{T}-u_{1}$, if $j=1$, then $\mathcal{T}-v_{1}$). Thus, from $\mathcal{D}^{m}(\mathcal{T})$, we can determine whether or not $i=1$.
Suppose $i=1$. Let $v_{1},\ldots,v_{m}$ be a maximal sequence of vertices satisfying that $v_{i}$ is matched to $u_{i}$ for all $1\leq i\leq m$. If $m=n$, all cards are identical and the tanglegram is reconstructable. Since m=n1m=n-1 is not possible, we consider m<n1m<n-1. Then, in exactly $m$ identical cards (corresponding to $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$), we have $v_{k}^{\prime}$ matched to $u_{k}^{\prime}$ for $1\leq k\leq m-1$ and $v_{m}^{\prime}$ not matched to $u_{m}^{\prime}$. In the remaining cards, $v_{k}^{\prime}$ and $u_{k}^{\prime}$ are matched for $1\leq k\leq m$. Thus, we can identify $m$, and the $m$ identical cards resulting from $\mathcal{T}-v_{i}u_{i}$ with $1\leq i\leq m$ allow us to reconstruct $\mathcal{T}$.
Suppose instead $i\neq 1$. In the cards of $\mathcal{D}^{m}(\mathcal{T})$, we consider which vertex is matched to $v_{1}^{\prime}$. From the cards other than $\mathcal{T}-v_{1}u_{i}$, we get i1i-1 cards with $v_{1}^{\prime}$ matched to $u_{i-1}^{\prime}$ (corresponding to $\mathcal{T}-u_{k}$ with $1\leq k<i$) and nin-i cards with $v_{1}^{\prime}$ matched to $u_{i}^{\prime}$ (corresponding to $\mathcal{T}-u_{k}$ with $i<k\leq n$). Which vertex $v_{1}^{\prime}$ is matched to on the remaining card, $\mathcal{T}-v_{1}u_{i}$, depends on the relationship of $i$ and $j$. The vertices matched to $v_{1}^{\prime}$ and their multiplicities in $\mathcal{D}^{m}(\mathcal{T})$ are collected in Table 2.
Table 2. For 1<i<n21<i<n-2, the vertices matched to $v_{1}^{\prime}$ in $\mathcal{D}^{m}(T)$ with multiplicities.
$j>i+1$ $j=i+1$ j=i1j=i-1 j<i1j<i-1
@@
$u_{i-1}^{\prime}\times(i-1)$ $u_{j}^{\prime}\times(1)$
From Table 2, if $v_{1}^{\prime}$ is matched to exactly two $u_{k}^{\prime}$, then $j=i+1$ or j=i1j=i-1 and the multiplicities allow us distinguish between these cases. If $j=i+1$, then there is a maximal sequence of vertices $v_{1},\ldots,v_{m}$ satisfying that $v_{k}$ is matched to $u_{i+k-1}$ for $1\leq k\leq m-1$. Hence, there are $m$ identical cards (corresponding to $\mathcal{T}-v_{i}$ with $1\leq i\leq m$) with $v_{k}^{\prime}$ matched to $u_{i+k-1}^{\prime}$ for $1\leq k\leq m-1$ and $v_{m}^{\prime}$ not matched to $u_{i+m-1}^{\prime}$. However, in all other cards, $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ are matched to a set of $u$’s that are consecutively and increasingly indexed. Thus, we can identify $m$, and one of the $m$ identical cards resulting from removing $v_{k}u_{i+k-1}$ with $1\leq k\leq m$ allows us to reconstruct $\mathcal{T}$. If instead j=i1j=i-1, then a similar argument with $v_{k}u_{i-k+1}\in\sigma$ for $1\leq k\leq m$ shows that $\mathcal{T}$ is reconstructable.
Finally, we consider when $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ for three different values of $k$. If these indices are non-consecutive, then we can easily identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$. In this case, $\mathcal{T}$ is reconstructable from this card. If the indices are consecutive, then $j=i+2$ or j=i2j=i-2 and we consider the multiplicity of the cards in which $v_{1}^{\prime}$ is matched to $u_{k}^{\prime}$ with $k$ maximum. In the former case, there is one such card and in the latter case, there are $n-i\geq 2$ such cards (since $i\leq n-1$). Hence, we can distinguish between these cases, identify both $i$ and the card $\mathcal{T}-v_{1}u_{i}$, and reconstruct $\mathcal{T}$. ∎
6. Caterpillar tanglegrams with a type 1 left or right tree
If the left or right tree of the caterpillar tanglegram is a type 1 tree, it is enough to consider the case when the left tree is $C_{n}$ and the right tree is type 1. We will use the distance-to-root labeling $\{u_{1},\ldots,u_{i}\}$ for the strippable leaves of $R$, and the corresponding distance to root labeling $\{u_{1}^{\prime}\ldots,u^{\prime}_{i^{\prime}}\}$ for the strippable leaves of the right subtree of a card $(C_{n-1},R^{\prime},\sigma^{\prime})$ whenever the right tree $R^{\prime}$ of the card is type 1.